# SoK: Runtime Integrity

### v2

@@ Abstract
This paper provides a systematic exploration of runtime integrity mechanisms, such as Control Flow Integrity (CFI) and Control Flow Attestation (CFA). It examines their differences and relationships while addressing crucial questions about the goals, assumptions, features, and design spaces. It includes examining a potential coexistence of CFI and CFA on the same platform. Through a comprehensive review of existing defenses, this paper positions CFI and CFA within the broader landscape of runtime defenses, critically evaluating their strengths, limitations, and trade-offs. The findings emphasize the importance of further research to bridge the gaps between CFI and CFA, advancing the field of runtime defenses.
@@ 1 Introduction
Unsafe programming languages like C and C++ are still prevalent, especially for lower-level system development [1]. Memory safety bugs, such as buffer overflows, are prominent enablers of attacks on programs written in such languages. Attacks that modify/inject code can be (to some extent) mitigated by existing defenses. Among them, Data Execution Prevention (DEP) and Write-Xor-eXecute (WX) [2] policies can prevent user-space code injection attempts at runtime. Secure boot can locally enforce boot-time code integrity (including the integrity of privileged software, e.g., stage 1 and 2 boot-loaders and kernel)[3, 4]. Static (i.e., boot-time) Remote Attestation (RA) [5, 6, 7] can further convince a remote party of the integrity of the booted code chain.
On the other hand, code-reuse attacks [8] (exemplified by Return-Oriented Programming – ROP [9] – and Jump-Oriented Programming – JOP [10]) need not modify the installed code, posing a significant threat even in the presence of the aforementioned defenses. They instead exploit memory corruption vulnerabilities to trigger out-of-order execution of sub-sequences of instructions (known as gadgets) within a program. This can result in unintended behavior even when code modifications are prevented. As illustrated in Figure 1, code-reuse attacks can be broadly classified into two categories: control flow hijacking and data-only attacks. The former directly corrupts memory storing code pointers, e.g., return addresses [9] and function pointers [10] during execution. The latter changes control flow related data, e.g., loop/conditional variables or counters, without causing control flow transfers that do not exist in the Control Flow Graph (CFG) of the target program [11, 12].
@@
In a parallel line of efforts, Control Flow Attestation (CFA) [34, 35, 36] has been proposed to enable remote Verifier(s) () to ascertain the execution integrity (including the absence of control flow attacks/violations) of an operation of interest performed by a remote device (called a prover or ). In its ideal form, CFA generates an authenticated log containing all dynamically defined control transfers occurring during the execution of an attested operation of interest111Note that some CFA variations aim at enabling continuous verification of all control flow transfers on , rather than focusing on individual operations of interest. For more details, see Section 2.3.. Nonetheless, similar to CFI, coarser-grained CFA approaches are also possible[37], establishing trade-offs between the accuracy of CFA evidence and performance, especially when overheads related to storage and transmission of said evidence to are a concern (e.g., when is a resource-constrained embedded platform).
Notably, not all CFA techniques guarantee that CFA evidence is received by . While this is sufficient for attestation, wherein a would not trust responses/values received from unless accompanied by CFA evidence, it does not support secure runtime auditing[38]. The latter aims to ensure that CFA evidence always reaches , even if is compromised, allowing for attack root cause analysis and appropriate remediation.
CFI and CFA goals can be viewed as runtime analogs of boot-time code integrity guarantees offered by secure boot vs. static RA. While CFI enables in loco detection of control flow violations (typically triggering exceptions when detected), CFA concerns providing remotely verifiable (unforgeable) evidence of the control flow path followed by an operation of interest executed by a device, thus enabling control flow path analysis by a remote .
@@ 1.1 Motivation & Intended Contributions
What are the assumptions, features, and design spaces of CFI vs. CFA, as well as similarities and differences?
@@
Could/should CFI and CFA coexist on the same platform?
Additionally, there is often confusion surrounding the terminology in the context of control flow-related mechanisms (e.g., prevention vs. local detection vs. remote detection; runtime attestation vs. runtime auditing; fine-grained vs. coarse-grained approaches; etc.) and their relationship to memory safety and compartmentalization defenses. This ambiguity makes it challenging to precisely understand the guarantees provided by each approach. Therefore, it becomes crucial to delve into such nuances to clearly grasp the benefits of each approach and their positions within the broader landscape of runtime software defenses.
@@ 1.2 Literature Selection Criteria
The selection criteria for the inclusion of academic papers or industrial proposals in our systematization are as follows:
@@ 1.3 Scope & Related Systematizations
Memory safety [39, 40, 41, 42] approaches aim to eliminate or reduce vulnerabilities that could lead to control/data flow attacks and data corruption during software development, i.e., before deployment. These typically work in two ways. First, memory safety can be a built-in security feature of programming languages such as Go and Rust. Rust [43], for instance, utilizes static compile-time analysis to optimize safety checks and memory management decisions, such as bounds check elimination while incorporating mechanisms (e.g., value ownership and borrowing) to ensure temporal safety. Second, memory safety can be obtained as memory-safe dialects of memory-safe programming languages. An example of this is Checked C [44], which augments C with spatial memory safety checks introduced at compilation time and runtime. This involves refining the C type system with safe pointer and array types with stricter usage models. Even so, this approach provides partial protection and presents compatibility challenges with legacy software.
A third class involves fortifying and isolating code with runtime checks [45], including compartmentalization [46], software fault isolation [47], and memory layout randomization [48]. While the focus here is on run-time violation detection and isolation, some texts still refer to this third class as memory safety techniques. In general, there is no consensus on whether the term should be restricted for the first two cases or include this third class (and more), leading to general confusion about whether the term “memory safety” should encompass [39].
@@
2 A Lighting Tour
This section reviews code-reuse attacks and existing defenses highlighting the role of CFI and CFA in this landscape.
@@ 2.1 Code Reuse Attack Background
Figure 1 illustrates the general classes of memory corruption-based attacks. Code-reuse attacks are further classified into control-flow hijacking and data-only attacks. At a high level, their difference lies in the former performing control flow transfers that do not exist in the legitimate CFG of the target program and the latter causing unintended transfers via edges that exist in the CFG. The two cases are depicted in Figure 2. Return Oriented Programming (ROP) [9] and Jump Oriented Programming (JOP) [10] are the two main categories of control-flow hijacking attacks. Both ROP and JOP stitch out-of-order sub-sequences of instructions, so-called gadgets, to modify the control flow path of the target program and perform a malicious behavior. As their names indicate, ROP corrupts backward edges, targeting gadgets that end with return instructions. JOP corrupts forward edges, targeting gadgets that end with indirect jump or call instructions.
Data-only attacks can be classified into two categories based on the type of non-control data manipulated [52]: Direct Data Manipulation (DDM) and Data-Oriented Programming (DOP). DDM attacks can be as simple as illegally modifying the value of a variable [11]. DOP attacks [12] on the other hand aim to perform expressive computation (often Turing-complete) allowing the by chaining carefully selected DOP gadgets so that the gadget chain exists as a valid path to the CFG. This is typically accomplished by corrupting control data, such as variables that define paths in conditional statements and loop counters.
@@ 2.2 Runtime Defenses
Figure 3 depicts the relationship between memory vulnerabilities, runtime exploits, and associated defences. The primary categories of runtime defenses against memory corruption-based attacks are illustrated in ( - ), which have been adapted from [14] and [45] respectively. Software testing tools, such as sanitizers [49] and fuzzers [53], act as a front-line defense in the pre-deployment phase, where the main goal is to find all possible vulnerabilities and fix them. Boot- and Load-time software verification mechanisms, such as Secure Boot [54], Measured boot [55], and Load-time attestation (e.g., the Linux Integrity Measurement Architecture (IMA) for user-space software [56]), are deployed as a primary shield in the post-deployment phase to prevent booting/loading of non-authentic software. However, the presence of memory corruption vulnerabilities at runtime remains a concern. Therefore, several runtime defenses have been proposed, each targeting specific steps of the attack process. Considering the five distinct attack steps ( - )) outlined in the general model of memory corruption attacks from [14], Figure 3 illustrates which class of defenses can counter each type of exploit and at which step. Note that, as information leaks are not integrity violations, they are not considered in the taxonomy shown in Figure 1. In the following, we summarize the individual attack steps and relevant defenses:
@@
Memory Vulnerability: Finding and exploiting a memory corruption vulnerability is an essential requirement for any of the attacks considered in Figure 3. Illegal access to a memory address, whether to read, write, or both, depends on the particular vulnerability. We note that vulnerabilities that enable read-only access are (by themselves) not sufficient to corrupt execution integrity of the target program.
@@
Integrity Violation: Exploiting vulnerabilities that grant illegal write access enables adversaries to tamper with the various aspects of a program, including the (i) program’s code (instructions in memory), (ii) control data (e.g., return addresses and function pointers), and (iii) non-control data (e.g., data variables and pointers). Isolation and compartmentalization mechanisms play a crucial role in enforcing access control permissions to mitigate integrity violations. These mechanisms restrict the targets that adversaries can access, often thwarting attacks at an early stage or preventing their spread to the rest of the system. For instance, access control mechanisms like the AArch64 (Un)Privileged Execution Never feature [57] make it significantly challenging to directly corrupt program code [58]. Code Pointer Integrity (CPI)[59] is a security mechanism that safeguards all code pointers and data pointers pointing to code by storing them in an isolated memory area. Code Pointer Separation (CPS)[59], a variant of CPI, isolates only code pointers while leaving the protection of data pointers to other measures for performance reasons. Software Fault Isolation (SFI) [47], memory tagging [60], and capability-based architectures (exemplified by CHERI [46]) operate at various granularity levels to isolate larger software components into distinct protection domains. These mechanisms limit the consequences of attacks that exploit memory vulnerabilities by confining them within specific compartments.
@@
Exploit Payload: If previous defenses are bypassed, adversary can inject payloads to manipulate the data and control flows of the target program. In general, the payload injection process requires knowledge of the program’s memory layout. In response, software diversification aims to impede the crafting of reusable exploits by introducing uncertainty through randomization. For instance, Address Space Layout Randomization (ASLR) [48] and Instruction Set Randomization (ISR) [61] are lightweight defenses that randomize memory layouts, making payload injection more challenging for control-flow hijacking and code-injection attacks. Additionally, Data Space Randomization (DSR) [62] can complicate data-only attacks. While these techniques offer probabilistic guarantees, they significantly raise the difficulty of runtime attacks.
@@
Exploit Dispatch: To successfully launch sophisticated attacks, adversary needs to divert the target program to operate on the injected payload. This step is crucial for expressive code-reuse attacks such as ROP, where the attack is initiated by manipulating the stack pointer to execute a sequence of selected gadgets in a predetermined order, with each gadget returning to a specific memory address in the following gadget to implement the desired attack behavior. Control Flow Integrity (CFI) [15] and Data Flow Integrity (DFI) [63] are two commonly used defenses to ideally detect and block control-flow hijacking and data-only attacks at this stage. These techniques involve implementing and enforcing policies that must be followed during program execution. However, contemporary literature shows that maintaining gap-free policies is inherently challenging, leaving potential exploit opportunities [64, 20, 65].
@@
Exploit Execution: As discussed above, ensuring the complete integrity of a victim program can be challenging. As a result, runtime attestation mechanisms have been proposed as a last line of defense to enable remote verification of code and execution integrity in a trustworthy manner. These mechanisms aim to detect tampering with code or violations of control/data flow. In addition, they also provide means to convince a remote party of the execution integrity of the target program during an operation of interest and enable in some cases auditing root cause vulnerabilities in case of exploits. In addition to measures such as [66] and DEP [2] policies, which are deployed to prevent code-injection attacks, remote attestation approaches of code binary [67, 7, 68, 69] are widely regarded as essential for providing remotely verifiable evidence of binary integrity at runtime. At any time during execution, they can be used to attest that the code (including instrumentation in it, e.g., as required by several CFI/CFA defenses) remains untampered. RA becomes paramount for most-privileged code (and single privilege systems, e.g, bare-metal micro-controllers) where full disablement of runtime code modifications implies the inability to perform remote software updates [70]. Control Flow Attestation (CFA) [34, 36] and Data Flow Attestation (DFA) [71, 72, 73] approaches have emerged to specifically detect and audit code-reuse attacks, enabling trustworthy remote verification of control-flow and data-flow integrity.
@@
As shown in Figure 3, attestation mechanisms build atop Roots of Trust (RoTs) as a foundation to provide trustworthy evidence of system/software state that can be remotely verified. For instance, RoTs are utilized in several key aspects of attestation mechanisms. They serve as a foundation for securely measuring system state and/or installed software (RoT for Measurement), securely storing attestation secret keys (RoT for Storage), and signing attestation reports (RoT for Reporting). Examples include Trusted Platform Modules (TPMs) [74], DICE [75], hardware extensions in Intel SGX-capable processors [76], and ARM TrustZone-based RoTs [77, 78], and various academic proposals, such as Keystone [79] and Byotee [80], among others.

### v3

@@ Abstract
This paper provides a systematic exploration of Control Flow Integrity (CFI) and Control Flow Attestation (CFA) mechanisms, examining their differences and relationships. It addresses crucial questions about the goals, assumptions, features, and design spaces of CFI and CFA, including their potential coexistence on the same platform. Through a comprehensive review of existing defenses, this paper positions CFI and CFA within the broader landscape of runtime defenses, critically evaluating their strengths, limitations, and trade-offs. The findings emphasize the importance of further research to bridge the gaps in CFI and CFA and thus advance the field of runtime defenses.
@@ 1 Introduction
Unsafe programming languages like C and C++ are still prevalent, especially for lower-level system development [1]. Memory safety bugs, such as buffer overflows, are prominent enablers of attacks on programs written in such languages. Attacks that modify/inject code can be (to some extent) mitigated by existing defenses. Among them, Data Execution Prevention (DEP) and Write-Xor-eXecute (WX) [2] policies can prevent user-space code injection attempts at runtime. Secure boot can locally enforce boot-time code integrity (including the integrity of privileged software, e.g., stage 1 and 2 boot-loaders and kernel)[3, 4]. Static (i.e., boot-time or load-time) Remote Attestation (RA) can further convince a remote party of the integrity of the booted code chain [5, 6].
On the other hand, code-reuse attacks [7] (exemplified by Return Oriented Programming – ROP [8] – and Jump Oriented Programming – JOP [9]) can pose significant threats without modifying the installed code, even in the presence of existing defenses. They instead exploit memory corruption vulnerabilities to trigger out-of-order execution of sub-sequences of instructions (known as gadgets) within a program. This can result in unintended behavior even when code modifications are prevented. As illustrated in Figure 1, code-reuse attacks can be broadly classified into two categories: control flow hijacking and data-only attacks. The former directly corrupts memory storing code pointers, e.g., return addresses [8] and function pointers [9] during execution. The latter changes control flow related data, e.g., loop/conditional variables or counters, without causing control flow transfers that do not exist in the Control Flow Graph (CFG) of the target program [10, 11].
@@
In a parallel line of efforts, Control Flow Attestation (CFA) [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] has been proposed to enable remote Verifier(s) () to ascertain the execution integrity (including the absence of control flow attacks/violations) of an operation of interest performed by a remote device (called a prover or ). In its ideal form, CFA generates an authenticated log containing all dynamically defined control transfers occurring during the execution of an attested operation of interest111Note that some CFA variations aim at enabling continuous verification of all control flow transfers on , rather than focusing on individual operations of interest. For more details, see Section 2.3.. Nonetheless, similar to CFI, coarser-grained CFA approaches are also possible [40], establishing trade-offs between the completeness of CFA evidence and performance, especially when overheads related to storage and transmission of said evidence to are a concern (e.g., when is a resource-constrained embedded platform).
Notably, not all CFA techniques guarantee that CFA evidence is received by . While this is sufficient for attestation, wherein a would not trust responses/values received from unless accompanied by CFA evidence, it does not support secure runtime auditing [41, 42]. The latter aims to ensure that CFA evidence always reaches , even if is compromised, allowing for attack root cause analysis and appropriate remediation.
CFI and CFA goals can be viewed as runtime analogs of boot-time code integrity guarantees offered by secure boot vs. static RA. While CFI enables in loco detection of control flow violations (typically triggering exceptions when detected), CFA provides remotely verifiable (unforgeable) evidence of the control flow path followed by an operation of interest executed by a device, thus enabling control flow path analysis by a remote .
@@ 1.1 Motivation & Intended Contributions
What are the assumptions, features, and design spaces of CFI vs. CFA, as well as their similarities and differences?
@@
Could CFI and CFA coexist on the same platform?
Additionally, there is often confusion surrounding the terminology in the context of control flow-related mechanisms (e.g., prevention vs. local detection vs. remote detection; runtime attestation vs. runtime auditing; fine-grained vs. coarse-grained approaches; etc.) and their relationship to memory safety and compartmentalization defenses. This ambiguity makes it challenging to precisely understand the guarantees provided by each approach. Therefore, it becomes crucial to delve into such nuances to clearly grasp the benefits of each approach and their roles within the broader landscape of runtime software defenses.
@@ 1.2 Literature Selection Criteria
The selection criteria for the inclusion of academic or industrial proposals in our systematization are as follows:
@@ 1.3 Scope & Related Systematizations
Memory safety [44, 45, 46, 47] approaches aim to eliminate or reduce vulnerabilities that could lead to control/data flow attacks and data corruption during software development, i.e., before deployment. These typically work in two ways. First, memory safety can be a built-in security feature of programming languages such as Go and Rust. Rust [48], for instance, utilizes static compile-time analysis to optimize safety checks and memory management decisions, such as bounds check elimination, while incorporating mechanisms (e.g., value ownership and borrowing) to ensure temporal safety. Second, memory safety can be obtained as memory-safe dialects of memory-safe programming languages. An example of this is Checked C [49], which augments C with spatial memory safety checks introduced at compilation time and/or runtime. This involves refining the C type system with safe pointer and array types with stricter usage models. Even so, this approach provides partial protection and presents compatibility challenges with legacy software.
A third category of techniques focuses on fortifying and isolating code through runtime checks [50]. These techniques include compartmentalization [51], software fault isolation [52], memory layout randomization [53], as well as CFI. While the primary goal of this class is to detect and isolate runtime violations, some literature still categorizes these methods as memory safety techniques. There is, however, no clear consensus on whether the term “memory safety” should be limited to the first two categories or expanded to include this third one (and possibly others). This lack of agreement has led to confusion over the scope of the term [44].
@@
2 A Lightning Tour
This section reviews code-reuse attacks and existing defenses, highlighting the role of CFI and CFA in this landscape.
@@ 2.1 Code Reuse Attack Background
Figure 1 illustrates the general classes of memory corruption-based attacks. Code-reuse attacks are further classified into control flow hijacking and data-only attacks. At a high level, their difference lies in the former performing control flow transfers that do not exist in the legitimate CFG of the target program and the latter causing unintended transfers via edges that exist in the CFG. The two cases are depicted in Figure 2. Return Oriented Programming (ROP) [8] and Jump Oriented Programming (JOP) [9] are the two main categories of control flow hijacking attacks. Both ROP and JOP stitch out-of-order sub-sequences of instructions, so-called gadgets, to modify the control flow path of the target program to perform a malicious action. As their names indicate, ROP corrupts backward edges, targeting gadgets that end with return instructions. JOP corrupts forward edges, targeting gadgets that end with indirect jump or call instructions.
Data-only attacks can be classified into Direct Data Manipulation (DDM) and Data Oriented Programming (DOP) based on the type of non-control data being manipulated [57]. DDM attacks can be as simple as illegally modifying the value of a variable [10]. DOP attacks [11], on the other hand, aim to perform expressive (often Turing-complete) computations by chaining carefully selected DOP gadgets, ensuring that the gadget chain forms a valid path within the CFG. This is typically achieved by corrupting non-control data, such as variables that define paths in conditional statements and loop counters.
@@ 2.2 Runtime Defenses
Figure 3 illustrates the relationship between memory vulnerabilities, runtime exploits, and associated defenses. The primary categories of runtime defenses against memory corruption-based attacks are illustrated in ( - ), which have been adapted from [13] and [50] respectively. Software testing tools, such as sanitizers [54] and fuzzers [58], act as a front-line defense in the pre-deployment phase, where the main goal is to find as many vulnerabilities as possible and fix them. Boot- and Load-time software verification mechanisms, such as Secure Boot [59], Measured Boot [60], and Load-time attestation (e.g., the Linux Integrity Measurement Architecture (IMA) for user-space software [61]), are deployed as a primary shield in the post-deployment phase to prevent booting/loading of non-authentic software. However, the presence of memory corruption vulnerabilities at runtime remains a concern after this stage. Therefore, several runtime defenses have been proposed, each targeting specific steps of the attack process. Considering the five distinct attack steps ( - ) outlined in the general model of memory corruption attacks from [13], Figure 3 illustrates which class of defenses can counter each type of exploit and at which step. As information leaks are not integrity violations, they are not considered in Figure 1. In the following, we summarize the individual attack steps and relevant defenses:
@@
Memory Vulnerability: Finding and exploiting a memory corruption vulnerability is an essential requirement for any of the runtime attacks considered in Figure 3. Illegal access to a memory address, whether to read, write, or both, depends on the particular vulnerability. We note that vulnerabilities that enable read-only access are (by themselves) not sufficient to corrupt the execution integrity of the target program.
@@
Integrity Violation: Exploiting vulnerabilities that grant illegal write access enables adversaries to tamper with the various aspects of a program, including the (i) program’s code (instructions in memory), (ii) control data (e.g., return addresses and function pointers), and (iii) non-control data (e.g., data variables and pointers). Isolation and compartmentalization mechanisms play a crucial role in enforcing access control permissions to mitigate integrity violations. These mechanisms restrict the targets that adversaries can access, often thwarting attacks at an early stage or preventing their spread to the rest of the system. For instance, access control mechanisms like the AArch64 (Un)Privileged Execution Never feature [62] make it significantly harder to directly corrupt program code [63]. Code Pointer Integrity (CPI)[64] is a security mechanism that safeguards all code pointers and data pointers pointing to code by storing them in an isolated memory area. Code Pointer Separation (CPS)[64], a variant of CPI, isolates only code pointers while leaving the protection of data pointers to other measures for performance reasons. Software Fault Isolation (SFI) [52], memory tagging [65], and capability-based architectures (exemplified by CHERI [51]) operate at various granularity levels to isolate larger software components into distinct protection domains. These mechanisms limit the consequences of attacks that exploit memory vulnerabilities by confining them within specific compartments.
@@
Exploit Payload: If previous defenses are bypassed, the adversary can inject payloads to manipulate the data and control flows of the target program. In general, the payload injection process requires knowledge of the program’s memory layout. In response, software diversification aims to impede the crafting of reusable exploits by introducing uncertainty through randomization. For instance, Address Space Layout Randomization (ASLR) [53] and Instruction Set Randomization (ISR) [66] are lightweight defenses that randomize memory layouts, making payload injection more challenging for control flow hijacking and code-injection attacks. Additionally, Data Space Randomization (DSR) [67] can complicate data-only attacks. While these techniques offer probabilistic guarantees, they significantly raise the difficulty of runtime attacks.
@@
Exploit Dispatch: To successfully launch sophisticated attacks, the adversary needs to divert the target program to operate on the injected payload. This step is crucial for expressive code-reuse attacks such as ROP, where the attack is initiated by manipulating the stack pointer to execute a sequence of selected gadgets in a predetermined order, with each gadget returning to a specific memory address in the following gadget to implement the desired attack behavior. Control Flow Integrity (CFI) [14] and Data Flow Integrity (DFI) [68] are two commonly used defenses to ideally detect and block control flow hijacking and data-only attacks at this stage. These techniques involve implementing and enforcing policies that must be followed during program execution. However, contemporary literature shows that maintaining gap-free policies is inherently challenging, leaving potential exploit opportunities [69, 19, 70].
@@
Exploit Execution: As discussed above, ensuring the complete integrity of a victim program can be challenging. As a result, runtime attestation mechanisms have been proposed as a last line of defense to enable remote verification of code and execution integrity in a trustworthy manner. These mechanisms aim to detect tampering with code or violations of control/data flow. In addition, they also provide means to convince a remote party of the execution integrity of the target program during an operation of interest and enable, in some cases, auditing root cause vulnerabilities in case of exploits. In addition to measures such as [71] and DEP [2] policies, which are deployed to prevent code-injection attacks, remote attestation approaches of code binary [72, 6, 73, 74] are widely regarded as essential for providing remotely verifiable evidence of binary integrity at runtime. At any time during execution, they can be used to attest that the code (including CFI/CFA instrumentation instructions) remains untampered. RA becomes paramount for most-privileged code (and single privilege systems, e.g. bare-metal micro-controllers) where full disablement of runtime code modifications implies the inability to perform remote software updates [75]. Control Flow Attestation (CFA) [33, 35] and Data Flow Attestation (DFA) [38, 37, 76] approaches have emerged to specifically detect and audit code-reuse attacks, enabling trustworthy remote verification of control and data flow integrity respectively.
@@
As shown in Figure 3, attestation mechanisms build atop Roots of Trust (RoTs) as a foundation to provide trustworthy evidence of system/software state that can be remotely verified. For instance, RoTs are utilized in several key aspects of attestation mechanisms. They serve as a foundation for securely measuring system state and/or installed software (RoT for Measurement), securely storing attestation secret keys (RoT for Storage), and/or signing attestation reports (RoT for Reporting). Examples include Trusted Platform Modules (TPMs) [77], DICE [78], hardware extensions in Intel SGX-capable processors [79], ARM TrustZone-based RoTs [80, 81], and various academic proposals such as Keystone [82] and BYOTee [83], among others.